OpenStudy (anonymous):

jamaican airways flies direct from miami international airport to montego bay.the plane flies for 1 hour 15 minutes to fly the 500 miles from miami to montego bay.what was the average air speed on the plane in miles per hour??

OpenStudy (anonymous):

500 miles /1.25 hours

OpenStudy (anonymous):

400 is the answer

OpenStudy (anonymous):

V=d/t V= 500/1.25 =400 mph

OpenStudy (anonymous):

ok i am correct